# Prospects for Observing High-redshift Radio-loud Quasars in the SKA Era: Paving the Way for 21-cm Forest Observations

### v1

@@ 3.4 Influence of Radio Spectral Index
Figures 8 and 9 show the influence of different radio spectral indices on the results. Typically, we use $-0.5$ as a standard to determine whether a radio spectrum is steep or flat. Early radio studies of galaxies suggested a steep radio spectral index of $-0.75$ (Condon, 1992), which has been adopted by multiple studies (e.g., Wang et al., 2007; Momjian et al., 2014; Bañados et al., 2015). We note that recently Gloudemans et al. (2021) used observational data from 93 quasars with $z>5$ to constrain the radio spectral index, obtaining a result of $-0.29$. Since our understanding of the radio spectra of quasars is still limited, here we calculate the influence of multiple values of radio spectral indices on the results.
Figure 8: Influence of different radio spectral index settings on the abundance of radio-loud quasars. Different colors of lines represent different spectral indices, while different line styles represent different radio flux thresholds (150 MHz flux). Sources with an abundance below the characteristic abundance marked by the red line are almost non-existent.
Figure 9: Influence of different radio spectral index settings on the forecast results of SKA-LOW observations. The observation frequency band of SKA-LOW is used for calculations here. Specifically, for continuum surveys, we used 150 MHz, while for 21-cm forest survey, we used the frequency of the redshifted 21-cm line.
3.5 Influence of Radio Loud Fraction
@@
Figures 10 and 11 illustrate the impact of the evolution of RLF with redshift and luminosity on the results. We adopted the evolutionary form from Jiang et al. (2007), which implies that the RLF at high redshift is much lower than 10%. Under the assumption of evolutionary RLF, the model predicts a significant decrease in the abundance of quasars observable by SKA-LOW’s one-year survey, making it more challenging to use the 21-cm forest method to probe cosmic reionization. If this RLF evolution is real, during the reionization period, the number of quasars with 100-hour integration times sufficient to resolve individual 21-cm lines is insufficient to support the 21-cm forest method, and we may need longer observation times to utilize relatively fainter quasars.
@@
Figure 10: Influence of evolving RLF on the abundance of radio-loud quasars. The blue line represents a constant 10% RLF, while the green line represents an evolving RLF. Different line styles indicate different radio flux thresholds (150 MHz flux). Sources with an abundance below the characteristic abundance marked by the red line are almost non-existent.
@@
4 conclusion
[…]
Another notable finding concerns the accretion behavior of AGNs. Our model, calibrated at high redshift, indicates that quasars have high accretion rates and short lifetimes. However, applying this setting to lower redshift results in significant discrepancies with observations. As a simple attempt, we recalibrated the model using the quasar luminosity function at $z=4$, resulting in an average Eddington ratio of $\lambda_{\rm Edd}=0.04$ and an average quasar lifetime of $t_{\rm q}=10^{8.2}$ years. This suggests that the accretion behaviors of early and late AGNs differ greatly.
The most significant result derived from our model is the prediction of the abundance of radio-loud quasars observable by SKA-LOW in a one-year survey. According to the standard parameter settings (Figure 7), we expect to identify over a thousand radio-loud quasars at $z>6$ that are sufficient to resolve individual 21-cm lines, within a year of observation. Even at higher redshifts, such as $z>9$, we can still detect approximately 20 such radio-loud quasars.
We tested the impact of various quasar radio spectral indices on our results (Figure 9). Our findings indicate that a steeper radio spectrum is more favorable for observations in the SKA-LOW frequency band. However, the radio spectral index has minimal effect on the results.
Our results suggest that the evolution of RLF is a looming issue for 21-cm forest in the SKA era (Figure 11). If the RLF remains close to 10% at high redshift, then the 21-cm forest can serve as an effective probe of reionization. However, if the RLF evolution claimed by Jiang et al. (2007) is real, detecting 21-cm forest background sources will be challenging, and we may need longer integration times to utilize relatively fainter quasars.
@@ acknowledgments
We thank K. Shimasaku for providing us with the compiled data of the halo–SMBH mass relation. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200, 2022SKA0110203, 2020SKA0110401, and 2020SKA0110100), the National Natural Science Foundation of China (Grants Nos. 11975072, 11875102, and 11835009), the National 111 Project (Grant No. B16009), the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-092), the National Key R&D Program of China (Grant No. 2022YFF0504300), and the science research grants from the China Manned Space Project (Grants Nos. CMS-CSST-2021-B01 and CMS-CSST-2021-A02). LOFAR data products were provided by the LOFAR Surveys Key Science project (LSKSP; https://lofar-surveys.org/) and were derived from observations with the International LOFAR Telescope (ILT). LOFAR (van Haarlem et al., 2013) is the Low Frequency Array designed and constructed by ASTRON. It has observing, data processing, and data storage facilities in several countries, which are owned by various parties (each with their own funding sources), and which are collectively operated by the ILT foundation under a joint scientific policy. The efforts of the LSKSP have benefited from funding from the European Research Council, NOVA, NWO, CNRS-INSU, the SURF Co-operative, the UK Science and Technology Funding Council and the Jülich Supercomputing Centre.

### v2

@@ 3.5 Influence of Radio Spectral Index
Figures 11 and 12 show the influence of different radio spectral indices on the results. Typically, we use $-0.5$ as a standard to determine whether a radio spectrum is steep or flat. Early radio studies of galaxies suggested a steep radio spectral index of $-0.75$ (Condon, 1992), which has been adopted by multiple studies (e.g., Wang et al., 2007; Momjian et al., 2014; Bañados et al., 2015). We used the stacking result of Gloudemans et al. (2021), $\alpha=-0.29$, as the standard parameter setting. However, it is noteworthy that the worse sensitivity of FIRST compared to LoTSS may naturally flatten the obtained spectral index. Since our understanding of the radio spectra of quasars is still limited, here we calculate the influence of multiple values of radio spectral indices on the results. As shown in Figures 11 and 12, our model indicates that a steeper spectral index is more beneficial for SKA-LOW’s observations of high-redshift radio-loud quasars. This advantage arises because the radio-loudness $R$ we are fitting is defined at $5$ GHz, meaning that our results directly reflect the flux density at this frequency. A steeper radio spectrum produces stronger flux for sources within SKA-LOW’s observation frequency range ($50-350$ MHz). The integral results suggest that adopting a steeper spectral index ($\alpha=-0.75$) instead of the flat spectral index proposed by Gloudemans et al. (2021) ($\alpha=-0.29$) could increase the number of observable quasars at $z\sim 9$ by approximately $1.3$ times (from $12$ to $28$). According to the requirements outlined by Shao et al. (2023) for background sources in the 21-cm forest (10 quasars at $z=9$), we anticipate sufficient background sources can be observed even with a flat spectral index. Consequently, the spectral index is not a critical parameter influencing the success of the 21-cm forest method.
3.6 Influence of Radio-Loud Fraction
@@
Figures 13 and 14 illustrate the impact of the evolution of RLF with redshift and luminosity on the results. We adopted the evolutionary form from Jiang et al. (2007), which implies that the RLF at high redshift is much lower than 10%. Under the assumption of evolutionary RLF, the model predicts a significant decrease in the abundance of quasars observable by SKA-LOW’s one-year survey, making it more challenging to use the 21-cm forest method to probe cosmic reionization. If this RLF evolution is real, during the reionization period, the number of quasars with 100-hour integration times sufficient to resolve individual 21-cm lines is insufficient to support the 21-cm forest method, and we may need longer observation times to utilize relatively fainter quasars.
@@
Figure 13: Influence of evolving RLF on the abundance of radio-loud quasars. The blue line represents a constant 10% RLF, while the green line represents an evolving RLF. Different line styles indicate different radio flux density thresholds (at 150 MHz). Sources with an abundance below the characteristic abundance marked by the red line are almost non-existent.
4 conclusion
One of the interesting results obtained from our model is the average lifetime $t_{\rm q}$ of high-redshift ($z\gtrsim 6$) quasars. Assuming Eddington accretion, the observed high-redshift quasar luminosity function requires that the average lifetime of quasars given by the model is $\sim 10^{5.5}$ years (see Section 2.4). Although this result is much smaller than early theoretical estimates ($\sim 10^{7}$ years), it is close to recent constraints using Lyman-$\rm\alpha$ near-zone data. Note that determining the average lifetime of quasars in the model requires a given halo–SMBH mass relationship, and currently there is large uncertainty in the estimation of high-redshift SMBH and halo masses based on quasar spectra. It is foreseeable that with the accumulation of high-redshift observational data, our model can provide a more reliable average lifetime of quasars, which will help us understand the growth process of SMBHs in the early universe.
@@
Another notable finding concerns the accretion behavior of AGNs. Our model, calibrated at high redshift, resulted in a short quasar lifetime under the assumption of Eddington accretion. However, applying this setting to lower redshift results in significant discrepancies with observations. Additionally, we found that simply changing the quasar lifetime does not help, indicating that the assumption of quasars accreting at the Eddington rate at low redshifts is wrong. As a simple attempt, we recalibrated the model using the quasar luminosity function at $z=4$ and allowed both $\lambda_{\rm Edd}$ and $t_{\rm q}$ to vary simultaneously, resulting in an average Eddington ratio of $\lambda_{\rm Edd}=0.04$ and an average quasar lifetime of $t_{\rm q}=10^{8.2}$ years. This suggests that the accretion behaviors of early and late AGNs differ greatly.
The most significant result derived from our model is the predicted abundance of radio-loud quasars observable by the SKA-LOW in an one-year survey. According to the standard parameter settings (Figure 8), we expect to identify over a thousand radio-loud quasars at $z>6$ that are sufficient to resolve individual 21-cm lines, within a year of observation. Even at higher redshifts, such as $z>9$, we can still detect approximately 20 such radio-loud quasars.
We tested the impact of various quasar radio spectral indices on our results (Figure 12). Our findings indicate that a steeper radio spectrum is more favorable for observations in the SKA-LOW frequency band. However, the radio spectral index is not a critical parameter influencing the success of the the 21-cm forest method.
Our results indicate that whether RLF evolves is crucial for the observation of the 21-cm forest in the SKA era (Figure 14). If the RLF remains close to 10% at high redshift, then the 21-cm forest can serve as an effective probe of reionization. However, if the RLF evolution claimed by Jiang et al. (2007) is real, then it will be challenging to find background sources for detecting individual 21-cm absorption lines, and we may need longer integration times to utilize relatively fainter quasars. In this case, a statistical measurement of the 21-cm forest would be vital in alleviating the requirement on the brightness of background sources (Mack & Wyithe, 2012; Thyagarajan, 2020; Shao et al., 2023).
@@ acknowledgments
We thank K. Shimasaku for providing us with the compiled data of the halo–SMBH mass relation. This work was supported by the National SKA Program of China (Grants Nos. 2020SKA0110401, 2022SKA0110200, 2022SKA0110203, and 2020SKA0110100), the National Natural Science Foundation of China (Grants Nos. 12473001, 11975072, 11875102, and 11835009), the National 111 Project (Grant No. B16009), the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-092), the National Key R&D Program of China (Grant No. 2022YFF0504300), and the science research grants from the China Manned Space Project (Grants Nos. CMS-CSST-2021-B01 and CMS-CSST-2021-A02). LOFAR data products were provided by the LOFAR Surveys Key Science project (LSKSP; https://lofar-surveys.org/) and were derived from observations with the International LOFAR Telescope (ILT). LOFAR (van Haarlem et al., 2013) is the Low Frequency Array designed and constructed by ASTRON. It has observing, data processing, and data storage facilities in several countries, which are owned by various parties (each with their own funding sources), and which are collectively operated by the ILT foundation under a joint scientific policy. The efforts of the LSKSP have benefited from funding from the European Research Council, NOVA, NWO, CNRS-INSU, the SURF Co-operative, the UK Science and Technology Funding Council and the Jülich Supercomputing Centre.